February 15
Attending: Eric and Jay from NERSC and users Andrei, Marjorie and Jeff P.
Cluster Utilization: Filled to capacity
Atlas plans: many students finishing and new ones are starting so there has been a lull in jobs.
Andrei is using cvmfs. It is working and should be put on all nodes including the debug node.
Discussed squid for db access on the hill - should it be moved to PDSF? Not critical for now...
Upcoming downtimes: eliza4 maintenance
Procurements/new hardware: Got the new eliza5 storage yesterday.
sl302 retirement is done.
torque/pbs testing: io resource not working with torque yet... migration in late March?
grid issues: will create atlasftr account; Eric to get ATLAS BNL account; Discussed whether PDSF will serve up ATLAS data - not completely clear at this point.
Jay will be on vacation Thursday-Monday


